OpenStudy (anonymous):

Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)).

OpenStudy (aakashsudhakar):

Let's just start with the raw expressions and simplify before plugging in given values. f(g(x)) = f(x^2 - 6x - 7) = (x^2 - 6x - 7) + 8 f(g(x)) = x^2 - 6x + 1 Now let's plug in our given value, x = 2. f(g(2)) = (2)^2 - 6(2) + 1 = 4 - 12 + 1 = -7 f(g(2)) = -7

OpenStudy (anonymous):

g(2)=4-12-7=-15 -15+8=7